OpenStudy (anonymous):

(FTOC) Evaluate the integral: (x-1)/x^(1/2) dx, [9,1] Answer: -40/3 Me: 12 (x-1)/x^(1/2) dx = (1/2)x^(3/2) - (1/2)x^(1/2) so: [(1/2)(9)^(3/2) - (1/2)(9)^(1/2)] - [(1/2)(1)^(3/2) - (1/2)(1)^(1/2) = [(27/2)-(3/2)] - [0] =12 I must be doing something wrong on doing the (1/2)x^(3/2) - (1/2)x^(1/2) but I can not figure it out.

OpenStudy (anonymous):

\[\int\limits_{1}^{9} \frac{ x-1 }{\sqrt{x} } dx\]

OpenStudy (anonymous):

Simplify it first: ( x - 1 )/ √x = x^½ - x ^(-½) ∫ ( x½ - x ^(-½) ) dx = ⅔ x^3/2 - 2x^½ [ 1, 9 ] = ⅔ [ 9^(3/2) - 1 ] - 2 [ 3 - 1 ] = ⅔ [ 27 - 1 ] - 4 = 52/ 3 - 12/3 = 40/3